SARALANCE:

How do you simplify x + 9 x + 3 ?

TheKayla:

x+9x+3=(1+9)x+3=10x+3

TheKayla:

Try that

SARALANCE:

OH THANK YOU SO MUCH

Alexis1415:

@thekayla next time try explaining how you got that and how you do it for the users thank you.